# Bitcoin Under Volatile Block Rewards: How Mempool Statistics Can Influence Bitcoin Mining

### v1

@@ Abstract.
As Bitcoin experiences more halving events, the protocol reward converges to zero, making transaction fees the primary source of miner rewards. This shift in Bitcoin’s incentivization mechanism, which introduces volatility into block rewards, could lead to the emergence of new security threats or intensify existing ones. Previous security analyses of Bitcoin have either considered a fixed block reward model or a highly simplified volatile model, overlooking the complexities of Bitcoin’s mempool behavior.
In this paper, we present a reinforcement learning-based tool designed to analyze mining strategies under a more realistic volatile model. Our tool uses the Asynchronous Advantage Actor-Critic (A3C) algorithm to derive near-optimal mining strategies while interacting with an environment that models the complexity of the Bitcoin mempool. This tool enables the analysis of adversarial mining strategies, such as selfish mining and undercutting, both before and after difficulty adjustments, providing insights into the effects of mining attacks in both the short and long term.
Our analysis reveals that Bitcoin users’ trend of offering higher fees to speed up the inclusion of their transactions in the chain can incentivize payoff-maximizing miners to deviate from the honest strategy. In the fixed reward model, a disincentive for the selfish mining attack is the initial loss period of at least two weeks, during which the attack is not profitable. However, our analysis shows that once the protocol reward diminishes to zero in the future, or even currently on days when transaction fees are comparable to the protocol reward, mining pools might be incentivized to abandon honest mining to gain an immediate profit.
Bitcoin, Selfish mining, Undercutting, Reinforcement learning, A3C
@@ 1. Introduction
As of September 2024, Bitcoin (Nakamoto, 2008) holds a market capitalization of nearly 1.18 trillion dollars (mar, 2024), surpassing the annual Gross Domestic Product (GDP) of over 90% of countries (gdp, 2024). While Bitcoin has progressed without major security attacks thus far, its steady growth over the past decade and a half does not guarantee future security, highlighting the need to address potential vulnerabilities. Protecting Bitcoin’s economy requires ensuring the resilience of its incentivization mechanism. If consensus nodes find their rewards insufficient to cover costs, they may abandon the chain, risking a halt in blockchain operations. Bitcoin’s incentivization comes from two sources: the protocol reward for mining a block and the transaction fee paid by users. In Bitcoin’s early years, the protocol reward significantly outweighed transaction fees. However, due to the halving mechanism, which reduces the protocol reward every four years, transaction fees are expected to match, and eventually replace, the protocol reward as the sole source of incentivization. In fact, there have already been periods, such as in December 2023 and April 20, 2024, when transaction fees equaled or even exceeded the protocol reward (exp, 2024).
@@
As the balance between transaction fees and the protocol reward changes over time, the incentivization dynamics of Bitcoin change as well, potentially leading to the emergence of new security threats or intensifying existing ones. One consequence of transitioning to a transaction-fee-driven era is that the total reward per block can no longer be considered fixed. Block rewards become volatile, as different blocks include varying transactions with different levels of transaction fees. This volatility in block rewards alters the profitability of various mining strategies. For example, under the fixed reward model, when only a single fork exists in the network, miners are incentivized to mine on the longest chain (Sapirshtein et al., 2017). Since the block reward is fixed, there is no benefit in taking the risk of orphaning previous blocks. However, in a volatile block reward model, a miner might find it more profitable to attempt to orphan a previously mined block with a significantly high transaction fee to steal those transactions (Bar-Zur et al., 2022; Carlsten et al., 2016). Additionally, a miner may choose to mine a secret fork of blocks to prevent others from building on that fork, allowing the miner to include all available transactions from the mempool in its secret fork. These examples highlight the need to revisit the security analyses conducted under the fixed reward model (Zur et al., 2020; Sapirshtein et al., 2017; Grunspan and Pérez-Marco, 2018) to better understand Bitcoin’s future security in the context of volatile block rewards.
@@
Some research has analyzed Bitcoin’s security under the volatile block reward model. Carlsten et al. (Carlsten et al., 2016) raised concerns about Bitcoin security as the protocol reward diminishes to zero, introducing the undercutting attack and showing how selfish mining could become more threatening in a transaction-fee-driven era. While the paper highlights Bitcoin’s instability under the volatile reward model, the analysis relies on simplified assumptions, such as unlimited block space, allowing blocks to collect all available transaction fees. Additionally, the analysis is based on a basic selfish mining strategy, overlooking the optimal strategy and other actions that adversarial miners might use in the volatile block reward model. To improve mining analysis in the transaction-fee setting, the authors in (Bar-Zur et al., 2022) introduced the WeRLman framework, which combines deep reinforcement learning with a Monte Carlo Tree Search (MCTS) tool, inspired by the AlphaGo Zero implementation (Silver et al., 2017). They assumed occasional whale transactions with significantly high fees that can result in volatility of block rewards. The study showed that as Bitcoin undergoes more halvings, the minimum hash-rate needed for profitable deviation from honest mining decreases. While the WeRLman paper made progress in analyzing Bitcoin security, its mempool implementation and treatment of transaction fees do not accurately reflect real-world conditions. In WeRLman, a block can include either no whale transactions or just one, resulting in rewards of either 1 or $1+\texttt{fee}$, where fee is the whale transaction fee. Therefore, the WeRLman framework simplifies the sophisticated mempool environment to only the count of whale transactions, all with the same fee. Moreover, WeRLman is designed to obtain a near-optimal strategy only after adjusting the mining difficulty. However, in practice, if a miner deviates from the honest strategy, it takes an epoch of at least two weeks for the difficulty to be adjusted, a relatively long period that WeRLman overlooks in its analysis.
In this paper, we enhance the analysis of mining strategies in Bitcoin by incorporating a more realistic volatile block reward model that obtains the near-optimal strategies both before and after the difficulty adjustment. We introduce two methods for evaluating Bitcoin security: the simplified method and the enhanced method. In the simplified method, we analyze historical Bitcoin block data over various periods to establish the relationship between the time required to mine a block, referred to as block generation time, and the total amount of transaction fees included in the block. This relationship is used to develop a reward function that approximates the behavior of the mempool under the volatile reward model. To evaluate mining strategies before difficulty adjustment within this simplified volatile reward framework, we employ a Markov Decision Process (MDP)-based tool. Because the MDP can handle a limited state space, our tool calculates a lower bound for miners’ profitability before a difficulty adjustment. This lower bound is particularly useful for determining the threshold mining parameters at which deviating from honest mining becomes profitable.
In the enhanced method, we improve both the implementation of mempool behavior and the training approach. Specifically, we apply the Asynchronous Advantage Actor-Critic (A3C) reinforcement learning algorithm (Mnih et al., 2016) to derive a near-optimal mining strategy. In this algorithm, multiple agents (representing a payoff-maximizing miner) interact with their own instance of the environment (the Bitcoin blockchain) to simultaneously train a global policy. To better model the Bitcoin mempool in the A3C environment, we analyze historical Bitcoin mempool data to estimate the transaction fee arrival patterns. In our implementation, transactions are categorized into different groups based on the fee paid per unit of weight. For each category, we estimate how the total transaction weight grows over time, incorporating these estimates into our environment. Our A3C implementation considers a larger set of actions and states compared to models designed for the fixed reward system, enabling it to obtain near-optimal policies both before and after difficulty adjustments under the volatile reward model.
Our analysis shows that when transaction rewards become comparable to protocol rewards—whether in the future with a significantly low protocol reward or currently on days with relatively high transaction fees—miners do not need to wait for infrequent whale transactions to be incentivized to deviate from the honest strategy. The inherent competition among transactions to be included in a block creates an opportunity for payoff-maximizing miners to earn more than their fair share of rewards by deviating from the honest strategy. As more time passes after the last block is mined on the network, new transactions tend to raise their fees to improve their chances of being included in the next block. In this situation, if a malicious miner removes other miners’ blocks from the canonical chain, it can increase the generation time of its own blocks. This extended block generation time allows the miner to include transactions with higher fees in its blocks.
@@
Another interesting observation is that under the volatile reward model, a deviant strategy, such as a selfish mining attack, can become profitable immediately after being launched without needing a mining difficulty adjustment. As discussed in the literature (Grunspan and Pérez-Marco, 2018; Negy et al., 2020; Grunspan and Pérez-Marco, 2023; Sarenche et al., 2024b), under the fixed reward model, a selfish miner incurs losses during the initial epoch of attack when the mining difficulty has not yet adjusted, the duration of which is at least two weeks. This initial loss period is considered one of the main reasons such attacks have not been observed in the Bitcoin network. However, under the volatile reward model, this loss period is shortened or may disappear entirely, making selfish mining a more significant threat in this era.
Figure 1. Immediate profitability of selfish mining in the transaction-fee era.
Once the protocol reward reaches zero, a threatening pattern that can emerge in the Bitcoin mempool occurs when the mempool is relatively empty, and blocks primarily collect transactions that arrive during their generation time. This mempool pattern, observable in Bitcoin’s mempool during September 2024, is susceptible to deviant strategies such as selfish mining and undercutting attacks. The graph depicted in Figure 1 represents a lower bound for the percentage increase in a miner’s time-averaged profit before a difficulty adjustment, on a day when the protocol reward is zero, and the same mempool pattern occurs as on September 5, 2024. This time-averaged profit percentage increase is calculated for a miner with a mining share equal to $1/3$, as a function of its communication capability, which denotes the ratio of miners receiving adversarial blocks earlier in the event of a block race. The normal communication capability is assumed to be $50\%$. However, as demonstrated in (Göbel et al., 2016), for a well-distributed mining pool, its connectivity ratio can increase substantially. As illustrated in Figure 1, the initial loss period of selfish mining can be entirely eliminated for malicious miners in the transaction-fee era.
@@
  • Discussion on how a volatile block reward model can affect selfish mining profitability, both before and after the difficulty adjustment, presented in Section 3.
  • Implementation of an MDP-based tool to calculate a lower bound on selfish mining profitability under a simplified mempool implementation, presented in Section 4.
  • Implementation of an A3C-based tool to obtain a near-optimal mining strategy under a realistic mempool model, presented in Section 5.
@@
2. System Model and Definitions
@@
A brief discussion of mining attacks that an adversary can conduct in our model, including selfish mining and undercutting, is presented in Appendix A. Depending on the context, we may refer to adversarial nodes as an attacker, a selfish miner, or an undercutter. Although the longest chain fork choice rule outputs a deterministic chain in most cases, situations may arise in which there are multiple forks with the same height in the view of a mining node (the same-height fork race), making the chain selection somewhat subjective. Based on how mining nodes select the winning chain in the presence of multiple same-height forks, the honest nodes are divided into $2$ sub-groups:
@@
According to the definition above, in a same-height fork race, petty-compliant nodes choose the fork with the lower total transaction fees included. The rationale behind this choice is that the fork with fewer transaction fees leaves more fees in the transaction pool, allowing the next block mined on this fork to include a higher amount of fees. Assuming all mining nodes share the same view of the transaction pool, in a same-height fork race, the winning chain for a petty-compliant node is the one whose tip block (the highest block) was mined earlier and, as a result, includes lower fees.
@@
The adversarial communication capability specifies the ratio of altruistic nodes that select the adversarial chain as the winning one in a same-height fork race. Let $\alpha$ denote the mining power share corresponding to the adversary, $\beta$ denote the ratio of petty-compliant mining nodes to the total honest mining nodes, and $\gamma$ denote the adversarial communication capability. This implies that the adversarial, petty-compliant, and altruistic mining shares are equal to $\alpha$, $\beta(1-\alpha)$, and $(1-\beta)(1-\alpha)$, respectively. Under the same-height fork race, the ratio $\gamma(1-\beta)(1-\alpha)$ of mining share (referring to the altruistic nodes that receive the adversarial chain earlier) always mines on top of the adversarial chain.
@@
3. Selfish Mining Profitability
This section provides an intuitive discussion on selfish mining profitability in the context of Bitcoin. We explain why selfish mining can be a profitable deviant strategy and explore the effect of the difficulty adjustment mechanism (DAM) on its profitability. We also discuss why selfish mining always incurs an initial period of loss in the fixed reward block model. Additionally, we examine how transitioning to the volatile block reward model can influence selfish mining profitability both before and after the difficulty adjustment mechanism. This discussion can aid in better understanding the results obtained through implementation in Sections 4 and 5.

### v2

@@ Abstract.
The security of Bitcoin protocols is deeply dependent on the incentives provided to miners, which come from a combination of block rewards and transaction fees. As Bitcoin experiences more halving events, the protocol reward converges to zero, making transaction fees the primary source of miner rewards. This shift in Bitcoin’s incentivization mechanism, which introduces volatility into block rewards, leads to the emergence of new security threats or intensifies existing ones. Previous security analyses of Bitcoin have either considered a fixed block reward model or a highly simplified volatile model, overlooking the complexities of Bitcoin’s mempool behavior.
This paper presents a reinforcement learning-based tool to develop mining strategies under a more realistic volatile model. We employ the Asynchronous Advantage Actor-Critic (A3C) algorithm, which efficiently handles dynamic environments, such as the Bitcoin mempool, to derive near-optimal mining strategies when interacting with an environment that models the complexity of the Bitcoin mempool. This tool enables the analysis of adversarial mining strategies, such as selfish mining and undercutting, both before and after difficulty adjustments, providing insights into the effects of mining attacks in both the short and long term.
We revisit the Bitcoin security threshold presented in the WeRLman paper (Bar-Zur et al., 2023) and demonstrate that the implicit predictability of valuable transaction arrivals in this model leads to an underestimation of the reported threshold. Additionally, we show that, while adversarial strategies like selfish mining under the fixed reward model incur an initial loss period of at least two weeks, the transition toward a transaction-fee era incentivizes mining pools to abandon honest mining for immediate profits. This incentive is expected to become more significant as the protocol reward approaches zero in the future.
Bitcoin, Selfish mining, Transaction fee, Undercutting, Reinforcement learning, A3C
@@ 1. Introduction
As of December 2024, Bitcoin (Nakamoto, 2008) holds a market capitalization of nearly 1.8 trillion dollars (mar, 2024), surpassing the annual Gross Domestic Product (GDP) of over 94% of countries (gdp, 2024). While Bitcoin has progressed without major security attacks thus far, its steady growth over the past decade and a half does not guarantee future security, highlighting the need to address potential vulnerabilities. Protecting Bitcoin’s economy requires ensuring the resilience of its incentivization mechanism. If consensus nodes find their rewards insufficient to cover costs, they may abandon the chain, risking a halt in blockchain operations. Bitcoin’s incentivization comes from two sources: the protocol reward for mining a block and the transaction fee paid by users. In Bitcoin’s early years, the protocol reward significantly outweighed transaction fees. However, due to the halving mechanism, which reduces the protocol reward every four years, transaction fees are expected to match and eventually replace the protocol reward as the sole source of incentivization. In fact, there have already been periods, such as in December 2023 and April 20, 2024, when transaction fees equaled or even exceeded the protocol reward (exp, 2024).
@@
As the balance between transaction fees and the protocol reward changes over time, the incentivization dynamics of Bitcoin change as well, potentially leading to the emergence of new security threats or intensifying existing ones. One consequence of transitioning to a transaction-fee-driven era is that the total reward per block can no longer be considered fixed. Block rewards become volatile, as the inclusion of transactions with varying fee levels causes fluctuations in the total reward per block. This volatility in block rewards alters the profitability of various mining strategies. For example, under the fixed reward model, once a miner receives a new block that extends the longest adopted fork in its view, the miner is incentivized to accept the new block and start mining on top of it (Sapirshtein et al., 2017). Since block rewards are fixed, there is no advantage in taking the risk of orphaning an already-mined block. However, in a volatile block reward model, a miner might find it more profitable to attempt to orphan a newly mined block with significantly high transaction fees to steal those transactions (Bar-Zur et al., 2023; Carlsten et al., 2016), a strategy referred to as undercutting. Additionally, a miner may engage in selfish mining by creating a private fork of blocks to take advantage of including all available transactions from the mempool in its private fork. These examples highlight the need to revisit the security analyses conducted under the fixed reward model (Zur et al., 2020; Sapirshtein et al., 2017; Grunspan and Pérez-Marco, 2018) to better understand Bitcoin’s future security in the context of volatile block rewards.
@@
Some research has analyzed Bitcoin’s security under the volatile block reward model in a simplified scenario. Carlsten et al. (Carlsten et al., 2016) raised concerns about Bitcoin security as the protocol reward diminishes to zero, introducing the undercutting attack and showing how selfish mining could become more threatening in a transaction-fee-driven era. While the paper highlights Bitcoin’s instability under the volatile reward model, the analysis relies on simplified assumptions, such as unlimited block space, allowing blocks to collect all available transaction fees. Additionally, the analysis is based on a basic selfish mining strategy, overlooking the optimal strategy and other actions that adversarial miners might use in the volatile block reward model. To improve mining analysis in the transaction-fee setting, the authors in (Bar-Zur et al., 2023) introduced the WeRLman framework, which combines deep reinforcement learning with a Monte Carlo Tree Search (MCTS) tool, inspired by the AlphaGo Zero implementation (Silver et al., 2017). They assumed occasional whale transactions with significantly high fees that can result in volatility of block rewards. The study showed that the minimum hash rate required for a profitable deviation from honest mining, referred to as the security threshold, decreases as Bitcoin undergoes more halvings. While the WeRLman paper made progress in analyzing Bitcoin security, its mempool implementation and treatment of transaction fees do not accurately reflect real-world conditions. In WeRLman, a block can include either a normal transaction or a whale transaction, resulting in blocks with only two distinct fee levels. This assumption simplifies the complex mempool environment to consider solely the count of whale transactions, all sharing the same fee. Moreover, WeRLman is designed to derive a near-optimal mining strategy only in a setting where the mining difficulty has been adjusted according to the adopted strategy. However, in practice, if a miner deviates from the honest strategy, it takes an epoch of at least two weeks for the difficulty to be adjusted, a relatively long period that WeRLman overlooks in its analysis.
In this paper, we enhance the analysis of mining strategies in Bitcoin by incorporating a more realistic volatile block reward model that obtains the near-optimal strategies both before and after the difficulty adjustment. We begin our analysis by revisiting the Bitcoin security threshold presented in the WeRLman paper (Bar-Zur et al., 2023). We demonstrate that, due to the design of the WeRLman environment, an adversary in this setting can predict the arrival of the next whale transaction before deciding on its action—an ability we refer to as predictive capability. We argue that the predictive capability strengthens the adversary compared to a real-world adversary, resulting in an underestimation of the security threshold. We modify the WeRLman environment to transform it into a non-predictable setting, where the adversary decides its action solely based on the mempool status at the moment of block generation, without any foresight regarding the types of subsequent transactions. The Bitcoin security threshold obtained under the non-predictable WeRLman model, while still lower than that in the fixed block reward model, is higher than what can be achieved under the original WeRLman environment.
We proceed by introducing a simplified model to analyze mining strategies, in which the volatility of block rewards arises from the varying generation times of blocks. We argue that block reward volatility is not solely attributed to the infrequent arrival of whale transactions. The inherent competition among transactions to be included in the chain quickly causes blocks with longer generation times to include higher transaction fees, resulting in block reward volatility even in the absence of whale transactions. To this end, we build our simplified model upon the observed positive correlation between block generation time and its fee. Compared to the WeRLman environment, our simplified environment includes a greater number of blocks with varying fee levels, allowing us to analyze more fine-grained strategies. We employ a Markov Decision Process (MDP)-based tool to analyze the profitability of mining strategies under our simplified volatile model. Due to the limited number of states in the simplified model, the MDP tool can accurately measure an upper bound for the security threshold. The results of our implementation show that the Bitcoin security threshold under our simplified model is lower than the security threshold obtained under the non-predictable WeRLman model.
@@
In the next step, we analyze the profitability of mining strategies before difficulty adjustment. An interesting observation under the volatile reward model is that a deviant strategy, such as a selfish mining attack, can become profitable immediately after launch without needing to wait for difficulty adjustment. As discussed in the literature (Grunspan and Pérez-Marco, 2018; Negy et al., 2020; Grunspan and Pérez-Marco, 2023; Sarenche et al., 2024a), under the fixed reward model assumption, a selfish miner incurs losses during the initial epoch of attack when the mining difficulty has not yet adjusted, the duration of which is at least two weeks. This initial loss period is one of the main reasons such attacks have not been observed in the Bitcoin network. However, under the volatile reward model, this loss period is shortened or may disappear entirely, making selfish mining a more significant threat in the transaction-fee era. Using our MDP-based tool within the simplified model, we analyze the Bitcoin security threshold prior to difficulty adjustment, which has not been explored in the WeRLman paper. Figure 1 illustrates the Bitcoin security threshold in the upcoming year as Bitcoin undergoes more halvings, based on different adversarial communication capabilities [^1] Communication capability is defined as the ratio of honest miners who receive the adversarial block first in the case of a block race.. Miners with a mining share above this threshold are not only incentivized to deviate from the honest strategy, but also do not face any initial loss. For a typical communication capability of $50\%$, when the protocol reward approaches 0, the Bitcoin security threshold reduces to $16\%$, which is within the range of current mining pool shares.
Figure 1. Bitcoin security threshold before DAM
Due to the limited state space in the simplified model and the simplifying assumptions made for implementation, the simplified model calculates a lower bound on mining strategy profitability. To conduct a more precise analysis of mining strategies under the volatile reward model, we propose an enhanced model that improves both the implementation of mempool behavior and the training approach. Specifically, we apply the Asynchronous Advantage Actor-Critic (A3C) reinforcement learning algorithm (Mnih et al., 2016) to derive a near-optimal mining strategy. In this algorithm, multiple agents (representing a payoff-maximizing miner) interact with their own instance of the environment (the Bitcoin blockchain) to train a global policy simultaneously. To better model the Bitcoin mempool in the A3C environment, we analyze historical Bitcoin mempool data to estimate the transaction fee arrival patterns. In our implementation, transactions are categorized into different groups based on the fee paid per unit of weight. We estimate how the total transaction weight grows over time for each category, incorporating these estimates into our environment. Our A3C implementation considers a larger set of actions and states compared to models designed for the fixed reward system, enabling it to obtain near-optimal policies both before and after difficulty adjustments under the volatile reward model. The undercut action is modeled as a continuous action in our A3C model, helping the adversary determine when and for how long to undercut the tip of the canonical chain.
@@
  • Revisiting the security threshold of difficulty-adjusted Bitcoin mining from the WeRLman paper (Bar-Zur et al., 2023), based on a real-world adversary without predictive capability (Section 3).
  • Introducing a simplified volatile reward model that includes a greater number of block reward levels and an MDP-based tool to achieve a more precise security threshold (Section 4).
  • Analyzing the security threshold of Bitcoin mining under the volatile model before difficulty adjustment to highlight the security threat regarding the immediate profitability of adversarial strategies (Section 5).
  • Implementing an A3C-based tool to derive a near-optimal mining strategy under a realistic mempool model (Section 6).
@@
2. Preliminaries
2.1. System Model and Definitions
@@
Depending on the context, we may refer to adversarial nodes as an attacker, a selfish miner, or an undercutter. In selfish mining, an attacker withholds a newly mined block instead of immediately broadcasting it to the network. In undercutting, the attacker attempts to orphan the tip of the canonical chain, forcing the transactions it contains back into the mempool. A brief discussion of mining attacks that an adversary can conduct in our model, including selfish mining and undercutting, is presented in Appendix A.
Although the longest chain fork choice rule outputs a deterministic chain in most cases, situations may arise in which there are multiple forks with the same height in the view of a mining node (the same-height fork race), making the chain selection somewhat subjective. Based on how mining nodes select the winning chain in the presence of multiple same-height forks, the honest nodes are divided into $2$ sub-groups:
@@
In a same-height fork race, petty-compliant nodes choose the fork with the lower total transaction fees included. The rationale behind this choice is that the fork with fewer transaction fees leaves more fees in the transaction pool, allowing the next block mined on this fork to include a higher amount of fees. Assuming all mining nodes share the same view of the transaction pool, in a same-height fork race, the winning chain for a petty-compliant node is the one whose tip block (the highest block) was mined earlier and, as a result, includes lower fees.
@@
The adversarial communication capability specifies the altruistic node ratio that selects the adversarial chain as the winning one in a same-height fork race.
@@
Definition 2.5 (Security threshold).
For a given security parameter $\epsilon$, the security threshold of a mining strategy $\pi$ is defined as the minimum mining share required for the time-averaged profit of strategy $\pi$ to exceed the time-averaged profit of the honest strategy by at least $\epsilon$. The security threshold within an environment is defined as the minimum mining share for which the profitability of the corresponding optimal strategy exceeds the time-averaged profit of the honest strategy by at least $\epsilon$.
2.2. Overview of Selfish Mining Profitability
In Bitcoin, a difficulty epoch is defined as the period during which 2016 blocks are appended to the canonical chain. At the end of each epoch, the Difficulty Adjustment Mechanism (DAM) recalibrates the difficulty target for the subsequent epoch based on the hash power estimated from the previous epoch. Adversarial behaviors that destroy a portion of network’s mining power can cause DAM to underestimate the active mining power. This underestimation lowers mining difficulty, potentially making such behaviors profitable. In the following, we briefly review the well-established fact that, under the fixed reward model, adversarial behaviors like selfish mining are unprofitable before difficulty adjustment (Grunspan and Pérez-Marco, 2018; Sarenche et al., 2024b; Negy et al., 2020). However, they can become profitable over the long term after applying DAM.